OpenStudy (anonymous):

Find the volume of a cube that has a total area of 96 sq. in.

OpenStudy (unklerhaukus):

\[SA_{\text{cube}}=6\times s^2 \]\[V_{\text{cube}}=s^3\]

OpenStudy (unklerhaukus):

first use \(SA_{\text{cube}}=96 \text{in}^2\) to find \(s\)

OpenStudy (anonymous):

is it 64 cubic units

OpenStudy (unklerhaukus):

\[96 \text{in}^2=6\times s^2\] \[s=\sqrt{\frac{96\text{in}^2}{6}} =\cdots\]

OpenStudy (anonymous):

96/6=16

OpenStudy (anonymous):

sqrt of 16 = 4

OpenStudy (anonymous):

4^3=64 :)